OpenStudy (anonymous):

Derivative of f(x)=sin(x)cos^(2) (pi-x). I got cos^(3)(pi-x)+sin^(3)(pi-x). Is that right?

OpenStudy (anonymous):

looks right, let me check it

OpenStudy (anonymous):

no not quite

OpenStudy (anonymous):

lets agree that \[\cos(\pi-x)=-\cos(x)\] to start

OpenStudy (anonymous):

yup I got that :)

OpenStudy (anonymous):

so what you are really looking at is \[-\sin(x)\cos^2(x)\]

OpenStudy (anonymous):

then i would get rid of the \(\pi-x\) not the other way around

OpenStudy (anonymous):

in fact perhaps your answer is right, but i would not use \(\pi-x\) as the input, just us \(x\) to make life easier

OpenStudy (anonymous):

you should get \[-\cos^3(x)+2\cos(x)\sin^2(x)\] if i am not mistaken